This is another tough one that I'm having trouble with.
1.) x / x - 1 + x = 4x - 3 / x - 1

Found 2 solutions by stanbon, RAY100:
Answer by stanbon(48564)

x / x - 1 + x = 4x - 3 / x - 1
-----
x/(x-1) + x = (4x-3)/(x-1)
multiply thru by x-1 to get:-
x + x(x-1) = 4x-3
x^2 = 4x-3
x^2 - 4x + 3 = 0
(x-3)(x-1) = 0
------
x = 1 or x = 3
-----------------
But x cannot be 1 in the original equaltion.
----------------
Only solution: x = 3
=============================
Cheers,
Stan H.

Answer by RAY100(1637)

check how we see the problem
x/(x-1)+x = 4x-3/(x-1)
mult thru by (x-1)
x +x(x-1)=4x(x-1)-3
distributing
x+x^2-x=4x^2-4x-3
simplying
x^2=4x^2-4x-3
0=3x^2-4x-3
like to factor but need quadratic formula
let a=3, b=(-4), c=(-3)
x= (-(-4) =/- sqrt ( (-4)^2 -4(3) (-3))/2(3)
x=(4+/- sq rt (16+36))/6
x=(4+/-(7.21))/6
x= 1.869, or -.5352 Answer
checking in original eqn both match ok